OpenStudy (studydino1201):

-35=7t A. -5 B. -1/5 C. 28 D. 42

OpenStudy (studydino1201):

@Will.H

OpenStudy (will.h):

You just have to get rid of 7 by division So therefore -35/7 = t -5= t

OpenStudy (pratyush5):

Can I know your doubt in this question ?

OpenStudy (studydino1201):

thanks -5 was the one i had first but i just making sure

OpenStudy (will.h):

Yes I know. If you wanna make sure you can substitute the answer you had in the original equation such that -35 = 7 * -5 -35= -35 Which sets the equation correctly which means that -5 is a correct solution